A car started from rest and gained a velocity of 60 km/h in 10 seconds. Calculate acceleration of the car
1
Expert's answer
2018-09-26T10:18:09-0400
Initial velocity:
v_i=0 m/s.
Final velocity:
v_f=60 km/h=60 1000/3600 m/s=60/3.6 m/s.
The acceleration of the car:
a=(v_f-v_i)/t=(60/3.6-0)/10=1.67 m/s^2 .